OpenStudy (anonymous):

The table shows the relationship between the number of cups of cereal and the number of cups of raisins in a cereal bar recipe: Number of cups of cereal 15 18 21 Number of cups of raisins 5 6 7 Adam plotted an equivalent ratio on the coordinate grid below: Which of these explains why the point Adam plotted is incorrect?

OpenStudy (anonymous):

If the number of cups of cereal is 60, then the number of cups of raisins will be 120. If the number of cups of raisins is 60, then the number of cups of cereal will be 120. If the number of cups of raisins is 60, then the number of cups of cereal will be 20. If the number of cups of cereal is 60, then the number of cups of raisins will be 20.

OpenStudy (anonymous):

Alright, first thing is to look at the graph you were given and to see if there is any correlation between the numbers.

OpenStudy (anonymous):

Okay, firstly. Number of cups of cereal 15 18 21 Number of cups of raisins 5 6 7 Here, it shows that you can find how many cups of cereal there was by multiplying the number of cups of raisins by 3, right?

OpenStudy (anonymous):

@Robert136

OpenStudy (anonymous):

@texaschic101 can u explain it how @Robert136 do it he give me da answer den he explain it to me so i can know it

OpenStudy (texaschic101):

15 cups of cereal to 5 cups of raisins = 60 cups of cereal to x cups of raisins 15/5 = 60/x -- cross multiply 15x = (60)(5) 15x = 300 x = 300/15 x = 20 so if there is 60 cups of cereal, then there is 20 cups of raisin